A girl pushes her little brother on his sled with
Answer #1

7) Given that F= 300 N & d= 750 meters cos\Theta=00

a) We know that work done(W)=(F\timesd cos\Theta)=(300\times750\times1)=(225000 Joules)=(225 KJ) (since cos00=1)

b) Energy spent on an object is transferred as a work done on that object.Hence enegy used = (225 KJ)

c) Here given that t= 25 sec, hence power =(Rate of doing work) = (W/t) = (225 KJ/25) =(9 KJ/sec)

d) Here 'd' remains same but t=10 sec, power = (W/t) = (225 KJ/10) =(22.5 KJ/sec)

8) Here m=75 kg and wall does not moved

Work is said to be done only if the object is moved to a distance(d) from respective initial position.

a) Here wall does not moved,hence work done on the wall by him=(0) Joules

b) But,Energy is spent to move there may exist internal energy but physically no energy.

c) Power =(W/t) = (0/t) = (0 Joules/sec)
Answer #2
W=Fd =(300N)(750m) =225,000 N. m. Or 225,000 Joules
Answer #3

SOLUTION :


7.


a.


Work done 

= Force * distance moved

= 300 * 750

= 225000 J 

= 225 kJ (ANSWER)


b.


Energy used

= Work done 

= 225000 J 

= 225 kJ(ANSWER)


c.


Power used 

= energy used per sec

= 225000 / 25

= 9000  W

= 9 kW (ANSWER).


d.


Power used 

= energy used per sec

= 225000 / 10

= 22500 W

= 22.5 kW (ANSWER).



8.


a.


There is no movement of the object,

So, 

Work done = 0  J


b.


Energy spent 

= work done 

= 0  J


(However impulse is created there = Force applied * time)


c.


Power used = 0 W

answered by: Tulsiram Garg